OpenStudy (anonymous):

Write a polynomial function of minimum degree with real coefficients whose zeros include those listed. Write the polynomial in standard form. 8, -14, and 3 + 9i

OpenStudy (anonymous):

these are the choices f(x) = x4 - 11x3 + 72x2 - 606x + 10,080 f(x) = x4 - 303x2 + 1212x - 10,080 f(x) = x4 - 11x3 - 72x2 + 606x - 10,080 f(x) = x4 - 58x2 + 1212x - 10,080

OpenStudy (amistre64):

subtract x from each root, and multiply the whole mess together. since they are suggesting a 4th degree poly, you might want to include the conjugate of that complex number

OpenStudy (amistre64):

8-x, -14-x, and 3 + 9i-x (8-x)(-14-x)(3+9i-x) produces a x^3 function that has complex issues ... if we include the conjugate we obtain: (8-x)(-14-x)(3+9i-x)(3-9i-x) which gets us to one of the options

OpenStudy (anonymous):

I got the first portion which is x^2+6x-112, but i can't seem to get the portion pertaining to 3+9i-x

OpenStudy (amistre64):

this makes the drudgery of pushing a pencil a little simpler, unless you need the practice of course http://www.wolframalpha.com/input/?i=%288-x%29%28-14-x%29%283%2B9i-x%29%283-9i-x%29

OpenStudy (amistre64):

3+9i-x 3-9i-x -------- 9+27i-3x -27i +81+9i x -3x -9i x +x^2 ------------------------ 9 -6x +81 +x^2 x^2 -6x +90 multiply that by the first one you found

OpenStudy (anonymous):

Thank you so much!

OpenStudy (amistre64):

youre welcome